OpenStudy (anonymous):

Can someone please help me? y-1 3/10=4 5/6

OpenStudy (xishem):

Is it this?\[y-1\frac{3}{10}=4\frac{5}{6}\]

OpenStudy (anonymous):

Yes, sorry I don't know how to type it properly.

OpenStudy (xishem):

That's ok. OK, so first we need to convert them to improper fractions:\[y-\frac{13}{10}=\frac{29}{6}\]Does that part make sense? Do you know how that works?

OpenStudy (anonymous):

Ok, I forgot where to start.

OpenStudy (anonymous):

What is the next step? Please?

OpenStudy (xishem):

Yep. You just want to move the fractions to the same side:\[y=\frac{29}{6}+\frac{13}{10}\]Do you know what to do next?

OpenStudy (anonymous):

Did it become addition because we brought 13/10 to the other side?

OpenStudy (xishem):

Yeah. You have to add (13/10) to both sides. So on the left side it cancels and on the right side you get +(13/10).

OpenStudy (anonymous):

Do we find common denominators?

OpenStudy (xishem):

Yep. Try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is the answer 6 4/30?

hero (hero):

Don't forget to simplify 4/30

OpenStudy (anonymous):

6 2/15?

hero (hero):

6 + 2/15

OpenStudy (anonymous):

Thank you Xishem and Hero, I really appreciate your help!

OpenStudy (xishem):

Cheers (:.

hero (hero):

The thing is though, if I were solving it, I'd solve it slightly differently. y - (1 + 3/10) = 4 + 5/6 y = (4 + 5/6) + (1 + 3/10) y = (4 + 1) + (5/6 + 3/10) y = 5 + (25/30 + 9/30) y = 5 + (25 + 9)/30 y = 5 + (30 + 4)/30 y = 5 + 30/30 + 4/30 y = 5 + 1 + 4/30 y = 6 + 4/30 y = 6 + 2/15